OpenStudy (anonymous):

What value should be added to the expression to create a perfect square? x2 – 20x A. 25 B. –25 C. 100 D. –100

OpenStudy (anonymous):

@shadowhuntergaming

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

david help him

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

a^2-2ab+b^2

Parth (parthkohli):

\[x^2 - 20x+ \cdots\]\[= (x)^2 - 2(10)(x) + \cdots\]Try to see how similar it is with something like\[(a-b)^2 = a^2 - 2ab + b^2\]

OpenStudy (nurali):

x^2-2(10)(x)+(10)^2

OpenStudy (nurali):

x^2-20x+100

OpenStudy (nurali):

so 100 is right one.

OpenStudy (anonymous):

thanks so much @Nurali

OpenStudy (nurali):

My Pleasure.